# Perceptions of the Fairness Impacts of Multiplicity in Machine Learning

### v1

@@ 5.1 Summary of responses
5.2 Evaluation of hypotheses
@@ 5.3 Exploratory analyses
E1 (Differences between procedural and distributional fairness)
While the set of 4 Likert items has high internal agreement for both the “pre” and “post” timestamps, the changes that occur with time are not distributed randomly across the four items. Overall, there is not a significant difference between “pre” and “post” fairness scores (H). But the first two Likert items, which ask about distributional fairness (whether a system will make good/fair decisions), see the scores decrease by an average of 0.122 points per item, while the second two Likert items, which ask about procedural fairness (the appropriateness of using the AI system), increase with time by an average of 0.131 points per item. This difference is statistically significant according to the Mann-Whitney U-test (W=74118, p=0.0001).
@@
6 Discussion
6.1 Procedural vs distributional fairness
Fairness scores did not change in aggregate after learning about multiplicity (H), but E1 reveals divergent behavior between distributional and procedural fairness-oriented questions. The existing literature suggests that multiplicity harms are primarily procedural [8, 23, 44], which contradicts lay stakeholders’ expressed preferences. We hypothesize that thinking about a model’s outcomes – and how there may not be an objectively “right” answer as evidenced by multiplicity – triggers critical thoughts towards the AI’s capabilities as a form of algorithmic aversion [27]. Specifically, the participants may see multiplicity as a “mistake,” thereby trusting algorithms less. This same effect may not occur for the procedural fairness questions, perhaps because participants feel that the AI system is already inscrutable (it is known that there is an informational component of fairness [20], and lay stakeholders rate models as more fair when they understand them better, e.g., via explanations [28, 72, 83, 101]). Instead, maybe we see an effect where respondents acclimate to the idea of using AI for the proposed purpose over the course of the survey, thereby giving it higher procedural fairness ratings at the later timestamp. This idea is supported by E2, as participants who take longer to complete the survey increase their final fairness rating more, on average.
Another fairness risk proposed by the literature is “systemic arbitrariness” [23], the idea that many actors using the same technology creates uniformly bad outcomes for some individuals who may have received more diverse results otherwise. It is possible that participants extrapolated systematic impacts from the scenario descriptions, however, we suspect this did not happen consciously – and indeed, none of the participants mentioned systemic impacts or repeated model encounters in the open-ended question.
6.2 Revealed preferences at odds with the literature
@@
There is a mismatch between the multiplicity solutions proposed by work in computer science and philosophy and the revealed preferences of lay stakeholders. In the CS and philosophy literature, scholars typically suggest using a more complex model (e.g., via ensembles) [22, 76] or randomization [23, 44]. Participants in the survey expressed comfort with fancy – choosing this option more often than would be expected by random chance – but strongly disliked random (H). This contradicts the established philosophical view that lotteries are fair in situations where a scarce resource must be allocated [12, 88]. A possible discrepancy between our setup and a run-of-the-mill lottery is the inclusion of machine learning in the decision-making process. Perhaps randomization via a lottery is familiar, but the addition of machine learning suggests that there must be some way to “learn” the best allocation, thereby making randomization too crude of a tool.
By far, the most preferred multiplicity resolution technique was human (H). Participants also suggested this option frequently in the open-ended question (E5). This approach has received some attention in the adversarial ML literature in the context of certifying robustness with an abstain option [19, 55], a goal that is related to ascertaining whether a particular decision may change given multiplicity. However, technical papers typically view the operational details of abstention as too trivial to consider, when in fact it is not straightforward. It is easy to say that a human decision maker will step in, but this may be costly, time-intensive, and can have undesirable effects on the overall fairness of a system [13, 34, 93].
Finally, we note a discrepancy between the de-facto ML practice and lay stakeholders’ preferences. In general, ML practitioners select a single good model from a set of reasonable models. Study participants do not like this approach: defaulting to the original model’s decision (i.e., ignore) is chosen in head-to-head matchups with other options less than half the time (H) and is infrequently suggested in the open-ended question (E5). This result indicates that the public does not view the machine learning status quo as optimal, rather, there should be intentionality and additional scrutiny when multiplicity occurs.
One possible multiplicity resolution technique that we did not consider is asking the decision subject, as was suggested by several respondents in the open-ended question (E5). Prior research has shown that decision subjects perceive AI as more fair when they have personal involvement in the review or contestation process [63], so this may be a promising option – albeit one with operationalization challenges.
6.3 Impacts of task choice
airness perceptions might depend on how the impact of multiplicity is described – e.g., by drawing explicit comparisons to the performance of a human expert, a factor we purposely kept constant to evaluate the role of task characteristics.
@@ 6.3 Impacts of task choice
6.4 Implications
There is no one-size-fits-all solution
Our most impactful finding is that ignore is generally not preferred: this finding indicates that standard ML practice needs to adapt to consider multiplicity, such as by increasing transparency when multiplicity impacts a decision. Likewise, participants expressed different preferences for multiplicity resolution techniques depending on the task, indicating that there is not a single multiplicity-resolution technique that is optimal in all situations. When multiplicity is present to a high extent, it may be challenging to find a solution that is fair, accurate, and scalable. (For instance, participants preferred human in many settings, but if 80% of predictions can vary with multiplicity, there may be very little time- or cost-saving if using ML to reduce human labor.) In these cases, it also may not be scientifically valid to use machine learning as has been argued for “predictive optimization” tasks [95].
Model developers should be able to justify their choices
Model developers should be able to defend the trade-offs inherent to their algorithms, including why one model was chosen over another when multiplicity exists. If this justification is not provided, and if the existence of other good models is revealed by model developers or a third-party, it may lower stakeholders’ perceptions of the fairness of the outcomes. Furthermore, the fact that overall fairness perceptions stay the same after learning about multiplicity (H) suggests that other factors pertaining to procedural fairness – e.g., transparency and accountability – may be more significant than multiplicity to final fairness perceptions. Developers focused on improving procedural fairness should note, though, that multiplicity in the model selection process may complicate transparency efforts by making it more difficult to provide justification for the specific model choice. Developers should be willing to be creative and listen to stakeholders: for instance, involving decision subjects directly may increase fairness perceptions, as suggested by the open-ended responses (E5) and existing work [63].
Need to reconcile views on multiplicity
We hypothesized that survey participants would view ML models as less fair after learning about multiplicity (H) since work in computer science and philosophy makes strong arguments that multiplicity impacts fairness (section 2.2). Since H was not supported by our data, either (1) the experts are incorrect about the fairness risks of multiplicity, (2) lay stakeholders do not see fairness risks of multiplicity due to low technical literacy, or (3) our study design was insufficient to capture stakeholders’ true views on multiplicity. We discuss (3) in section 6.5, so we will overview (1) and (2) here.
If (1), the fairness risks of multiplicity summarized in section 2.2 must paint an incomplete picture. For instance, maybe multiplicity poses a theoretical fairness risk, but one that is is either too small to matter in practice, or that is overpowered by the benefits of using ML. Or, perhaps, the philosophical view that only systematic arbitrariness is a fairness risk [23] is correct. Articulating the fairness risks of multiplicity relative to other fairness issues in ML may help scholars clarify when, exactly, multiplicity poses a fairness risk.
If (2), then H was not supported because the stakeholders’ technical literacy was not high enough to realize that multiplicity poses the procedural fairness risks suggested by computer scientists and philosophers. As ML technology powers more and more decisions that used to be made by humans, it is important that stakeholders are able to recognize its limitations to avoid inappropriate reliance on and trust in ML systems [61]. One way that reliance and trust can be calibrated is through improving technical/AI literacy [16]. We discuss how our survey design could be modified to promote better understanding in section 6.5. But also, technical literacy can be built gradually over time by exposure to news stories and public discourse [69]. We encourage people working at the intersection of technology and communication to discuss multiplicity and to take a critical lens, more broadly, when describing new ML systems in socially-impactful domains.
6.5 Limitations and opportunities for future work
@@
A limitation of this work is that multiplicity is not an easy concept to grasp, especially for people without machine learning experience. We attempted to explain it clearly and in simple terms and iterated on the survey text through multiple pilot iterations. We attempted to check that participants understood multiplicity through an initial comprehension check, but some participants may have gotten lucky and passed this question by chance, only looked for the correct answer rather than engaging critically with the reading, or may have forgotten details when answering the remaining survey questions.
Unknown role of other factors
To counteract the cognitive load required to reason about multiplicity, we tried to make the other details about the tasks and models as simple as possible, including by providing only high-level details of the AI system’s functionality and context. Some participants may have felt like they did not have enough information to evaluate the system’s fairness, and the presence of multiplicity may have seemed insignificant relative to other unknown factors. In particular, we did not include any explanations of how the models made their decisions to keep the setup simpler. Informational transparency is a component of overall fairness [20], though, so future work could explore how different explanations of both model behavior and of multiplicity itself impact fairness perceptions. Likewise, we do not include personal impact – i.e., whether the prediction a model makes on a specific decision subject varies under multiplicity – as a factor due to the experiment’s complexity. We hypothesize – and future work should check – that if participants are told that the decision the model makes for them varies under multiplicity, fairness perceptions would decrease, mirroring work on outcome favorability [96]. To combat issues with the large number of unknowns inherent to hypothetical situations, future research could be situated in a real-world task, where explaining more details about the model would not necessitate a large cognitive load due to baseline familiarity with an existing algorithm. Existing research also shows that people behave differently in hypothetical settings [32], again highlighting the need for additional research situated in a real use case.
@@
We conducted our survey with US-based crowdworkers. It is known that fairness perceptions vary cross-culturally [5, 50], as do attitudes towards uncertainty [41]. It therefore seems likely that attitudes towards the fairness of multiplicity in ML would vary across cultures, too, which future work could explore.
7 Conclusions
@@
We conducted a study (n=357) of crowdworkers to determine how lay stakeholders (1) perceive the fairness of ML models when multiplicity impacts decisions, and (2) prefer that this multiplicity be resolved. Our results indicate that stakeholders have strong preferences about how multiplicity should be resolved. In particular, they express a preference for human decision makers, and a dispreference for both ignoring multiplicity (arbitrarily using a single model) and randomizing between different models’ outputs. These findings counter not only the established ML practice of looking for a single well-performing model, but also philosophical arguments that randomization is the fairest way to deal with multiplicity. Despite these strong opinions about resolving multiplicity, the presence of multiplicity did not impact how fair the participants thought the models were, which is a discrepancy that future work will need to address. One promising avenue is to consider how perceptions of informational fairness relating to multiplicity, e.g., the use of explanations, intersects with the procedural and distributional sides of fairness that we investigated. Ultimately, a greater understanding of when ML design choices – including how multiplicity is resolved – are perceived as fair can help guide model developers to create technology that aligns with stakeholders’ values.
Acknowledgements

### v2

@@ 5. Data Analysis
5.1. Summary of responses
5.2. Evaluation of hypotheses
@@ 5.3. Exploratory analyses
E1 (Differences between different dimensions of fairness)
While the set of 4 Likert items has high internal agreement for both the “pre” and “post” timestamps, the changes that occur with time are not distributed randomly across the four items. Overall, there is not a significant difference between “pre” and “post” fairness scores (H). But the first two Likert items, which ask about model outcomes (whether a system will be able to make good/fair decisions), see the scores decrease by an average of 0.122 points per item, while the second two Likert items, which ask about the appropriateness of using the AI system, increase with time by an average of 0.131 points per item. This difference is statistically significant according to the Mann-Whitney U-test (W=74118, p=0.0001).
The plot shows the marginal mean estimates for each multiplicity-resolution technique. We will report the lower bound, mean estimate, and upper bound for each. Default bad: 0.276, 0.328, 0.380. Default good: 0.522, 0.576, 0.630. Fancy: 0.551, 0.602, 0.653. Ignore: 0.367, 0.41, 0.464. Human: 0.749, 0.790, 0.832. Random: 0.235, 0.279, 0.323.
@@
6. Discussion
6.1. The standard ML approach is unfair (ignore is disliked)
Our most impactful finding is that ignore is generally not preferred.
Implication: Standard ML workflows need to adapt to consider multiplicity, such as by increasing transparency when multiplicity impacts a decision.
Depending on the target audience, interventions to increase transparency could take different forms. For instance, common machine-learning packages could alert developers when multiplicity exists (e.g., by automatically performing a bootstrapping analysis in the background). Model documentation should include an analysis of what proportion of test samples are subject to receive a conflicting prediction under multiplicity, so that anyone who decides to use or adapt the model in the future is aware. In public communications about a model, the fraction of inputs (based on validation data) that are subject to receive a different prediction under multiplicity could be reported.
Implication: Model developers should be able to justify how multiplicity is handled, whether this is by deferring to a human decision-maker, employing randomization, using ensemble methods, or something else.
It is possible that decision subjects would be okay with using a single model, and knowingly ignoring multiplicity, if this decision is justified. But, just as model developers need to justify decisions such as feature selection and statistical fairness trade-offs, they should justify why the existence of multiplicity does not pose an existential threat to their project. In particular, they should consider potential harms to individuals, e.g., whether the model’s use will be widespread and lead to systemic exclusion of some people.
6.2. Is multiplicity an inherent fairness risk?
Our results show that participants’ fairness perceptions did not decrease when multiplicity was present (i.e., H was not supported). This contradicts work in ML and philosophy that assumes and argues that multiplicity poses a fairness risk (see section 2.2). Since H was not supported by our data, either (1) the experts are incorrect about the fairness risks of multiplicity, (2) lay stakeholders do not see fairness risks of multiplicity due to low technical literacy, or (3) our study design was insufficient to capture stakeholders’ true views on multiplicity. We discuss (3) when describing our work’s limitations (section 6.6), so we will overview (1) and (2) here.
If (1), the fairness risks of multiplicity summarized in section 2.2 must paint an incomplete picture. Perhaps the philosophical view that only systematic arbitrariness is a fairness risk (Creel and Hellman, 2022) is correct. Alternatively, maybe multiplicity poses a theoretical fairness risk, but one that is is either too small to matter in practice, or that is overpowered by the benefits of using ML. Participants may feel as if the ML system is a closed box, and thus that its decisions are impenetrable. This perception would, understandably, lead participants to feel disenfranchised in their interactions with the ML model. From this perspective, multiplicity is not the primary reason why a model’s usage could be unfair.
Implication: Theoretical researchers should articulate how the fairness risks of multiplicity relate to other fairness issues in ML – such as accountability and recourse – to help clarify when, exactly, multiplicity poses a fairness risk. Future research about the perceived fairness of multiplicity should do more to access the “why:” what are the driving forces in fairness perceptions?
If (2), then H was not supported because the stakeholders’ technical literacy was not high enough to realize that multiplicity poses the procedural fairness risks suggested by computer scientists and philosophers. As ML technology powers more and more decisions that used to be made by humans, it is important that stakeholders are able to recognize its limitations to avoid inappropriate reliance on and trust in ML systems (Long and Magerko, 2020). One way that reliance and trust can be calibrated is through improving technical/AI literacy (Chiang and Yin, 2022). We discuss how our survey design could be modified to promote better understanding in section 6.6. But also, technical literacy can be built gradually over time by exposure to news stories and public discourse (Nguyen and Hekman, 2024).
Implication: We encourage people working at the intersection of technology and communication to discuss multiplicity and to take a critical lens, more broadly, when describing new ML systems in socially-impactful domains.
6.3. Randomization is disliked
@@
Scholars in ML and philosophy typically suggest using a more complex model (e.g., via ensembles) (Cooper et al., 2024; Roth et al., 2023) or randomization (Creel and Hellman, 2022; Jain et al., 2024). Participants in the survey expressed comfort with fancy – choosing this option more often than would be expected by random chance – but strongly disliked random (H). This contradicts the established philosophical view that lotteries are fair in situations where a scarce resource must be allocated (Broome, 1990; Sher, 1980). A possible discrepancy between our setup and a run-of-the-mill lottery is the inclusion of machine learning in the decision-making process. Perhaps randomization via a lottery is familiar, but the addition of machine learning suggests that there must be some way to “learn” the best allocation, thereby making randomization too crude of a tool.
Implication: More research should be conducted into why random is disliked, even though lotteries are generally viewed as fair. Researchers can investigate whether providing additional education about machine learning (e.g., how models can make mistakes and that there is often not a single, optimal solution) improves comfort with random.
6.4. Participants prefer human intervention, but is that scalable?
Participants generally prefer human over other multiplicity-resolution techniques (H). However, when large proportions of predictions can vary under multiplicity, deferring to a human decision-maker is difficult to operationalize and negates cost and time savings associated with using ML. In these cases, one option is to simply forgo using ML at all to avoid the appearance of automation and the expense of developing a system that only handles a small number of cases. (Indeed, it may not be scientifically valid to use machine learning when many predictions vary under multiplicity, as has been argued for “predictive optimization” tasks (Wang et al., 2024).) Instead, decision-makers could rely on an interpretable, publicly-available set of guidelines. Alternatively, model developers could use a strategy that combines fancy and human. Recall that fancy (e.g., using ensemble methods) is favored by ML researchers as a technique to reduce multiplicity. Model developers could use ensembles or more complex model architectures to reduce multiplicity, and then defer to a human decision maker only for the smaller number of cases whose predictions remain ambiguous. In our experiments, participants were generally comfortable with fancy, so while further research is needed to confirm that a combination fancy and human approach would be viewed as fair, we are optimistic about that possibility.
Implication: Model developers may not be able to align with stakeholders’ preferences for human in all cases, but should consider whether other, more interpretable decision-making rules are more appropriate when multiplicity is present. Alternatively, using ensemble methods to reduce multiplicity, and then deferring to a human on the smaller number of ambiguous cases, could strike a good balance.
6.5. Impact of task choice
@@
Implication: Model developers should involve participants in the design process to figure out how they prefer that multiplicity be handled for the specific task.
6.6. Limitations and opportunities for future work
@@
A limitation of this work is that multiplicity is not an easy concept to grasp, especially for people without machine learning experience. We attempted to explain it clearly and in simple terms and iterated on the survey text through multiple pilot iterations. Our results do not show significant differences in the responses for participants with or without ML experience. We attempted to check that participants understood multiplicity through an initial comprehension check, but some participants may have gotten lucky and passed this question by chance, only looked for the correct answer rather than engaging critically with the reading, or may have forgotten details when answering the remaining survey questions.
Unknown role of other task-related factors
To counteract the cognitive load required to reason about multiplicity, we tried to make the other details about the tasks and models as simple as possible, including by providing only high-level details of the AI system’s functionality and context. Although we tried to select tasks that are broadly relevant, some participants may not have seen the relevance to their own lives or imagined themselves as the decision subjects. Some participants may have felt like they did not have enough information to evaluate the system’s fairness, and the presence of multiplicity may have seemed insignificant relative to other unknown factors. In particular, we did not include any explanations of how the models made their decisions to keep the setup simpler. Informational transparency is a component of overall fairness (Colquitt, 2001), though, so future work could explore how different explanations of both model behavior and of multiplicity itself impact fairness perceptions. Likewise, we do not include personal impact – i.e., whether the prediction a model makes on a specific decision subject varies under multiplicity – as a factor due to the experiment’s complexity. We hypothesize – and future work should check – that if participants are told that the decision the model makes for them varies under multiplicity, fairness perceptions would decrease, mirroring work on outcome favorability (Wang et al., 2020). To combat issues with the large number of unknowns inherent to hypothetical situations, future research could be situated in a real-world task, where explaining more details about the model would not necessitate a large cognitive load due to baseline familiarity with an existing algorithm. Existing research also shows that people behave differently in hypothetical settings (FeldmanHall et al., 2012), again highlighting the need for additional research situated in a real use case.
Unknown role of other personal factors
We did not find significant trends between demographic groups for any of the factors we recorded. However, our analysis was not powerful enough to detect how any intersectional identities (e.g., gender and education level) may impact perceptions of multiplicity. Furthermore, our collection of some demographic attributes, especially experience with AI, may have been too coarse to see meaningful trends. Future work could consider contrasting specific participant groups (e.g., ML engineers versus adults with no ML experience) to gain a better sense of how fairness perceptions vary with ML experience.
@@
We conducted our survey with US-based crowdworkers. It is known that fairness perceptions vary cross-culturally (Berman et al., 1985; Kim and Leung, 2007), as do attitudes towards uncertainty (Hofstede, 1984). It therefore seems likely that attitudes towards the fairness of multiplicity in ML would vary across cultures, too, which future work could explore.
7. Conclusions
@@
We conducted a study (n=357) of crowdworkers to determine how lay stakeholders (1) perceive the fairness of ML models when multiplicity impacts decisions, and (2) prefer that this multiplicity be resolved. Our results indicate that stakeholders have strong preferences about how multiplicity should be resolved. In particular, they express a preference for human decision makers, and a dispreference for both ignoring multiplicity (arbitrarily using a single model) and randomizing between different models’ outputs. These findings counter not only the established ML practice of looking for a single well-performing model, but also philosophical arguments that randomization is the fairest way to deal with multiplicity. Despite these strong opinions about resolving multiplicity, the presence of multiplicity did not impact how fair the participants thought the models were, which is a discrepancy that future work will need to address. One promising avenue is to consider how perceptions of informational fairness relating to multiplicity, e.g., the use of explanations, intersects with the procedural components of fairness that we investigated. Ultimately, a greater understanding of when ML design choices – including how multiplicity is resolved – are perceived as fair can help guide model developers to create technology that aligns with stakeholders’ values.
Acknowledgements.